OpenStudy (anonymous):

Simplify. –6 ÷ (–2) – 6 + 4 help me

OpenStudy (dahlindubs11):

answer is 1... -6/-2=3... 3-6=-3... -3+4=1

OpenStudy (anonymous):

okay thank you so much can u help me on more?

OpenStudy (anonymous):

Simplify. [2.2 ÷ (4 – 2)] • (–3)

OpenStudy (anonymous):

someone help o-o

OpenStudy (dahlindubs11):

do the parentheses in the brackets first... 4-2=2... 2.2/2=1.1... 1.1*-3=-3.3

OpenStudy (anonymous):

ya i do that and sometimes i dont even get the right answer but thanks again

OpenStudy (dahlindubs11):

yeah always do brackets first. then parentheses, if necessary, in the brackets the go from there

OpenStudy (anonymous):

okay thanks i have like 2 more

OpenStudy (anonymous):

well one now

OpenStudy (dahlindubs11):

go ahead... can u pm me. cuz its all lagging